AshleeM:

A collection of nickels, dimes, and quarters totals $8.20. The number of nickels and dimes together is twice the number of quarters. The value of the nickels is one-third the value of the dimes. How many of each kind of coin are in the collection?

mhanifa:

You should build a system of equations to solve this. Consider nickel = n, dime = d, quarter = q. We know nickel = 5 cents, dime = 10 cents, quarter = 25 cents and $8.20 = 820 cents. The first equation would be: 5n + 10d + 25q = 820 The second equation: (n + d)*2 = q The third equation: n = 1/3d Let me know if further help is required.